A 114 gram tennis ball strikes the court ground vertically with a speed of 3 m/s. If the ball bounces back with a speed of 2.8 m/s vertically upward, find the magnitude of the change in the tennis ball’s momentum, in kg m/s.
Answer #1

mass of the tennis ball, m = 114 gm = 0.114kg

the initial velocity of the ball, u = -3 m/s ( upward direction is assumed positive)

the final velocity of the ball, v = 2.8 m/s

magnitude of the change in momentum of the tennis ball, p = m ( v - u) = 0.114* ( 2.8 - (-3)) = 0.6612 kg.m/s (Ans)